Inverse Laplace Transforms

Find L^-1 {F(s)} when F(s) is given by


3. s-1/s^2 (s+3)


1
Expert's answer
2022-05-26T11:38:12-0400
"\dfrac{s-1}{s^2(s+3)}=\dfrac{A}{s+3}+\dfrac{B}{s}+\dfrac{C}{s^2}""=\dfrac{As^2+Bs^2+3Bs+Cs+3C}{(s+3)(s+7)}""A+B=0""3B+C=1""3C=-1"


"A=-\dfrac{4}{9}, B=\dfrac{4}{9}, C=-\dfrac{1}{3}"




"L^{-1}(\dfrac{1}{(s+3)(s+3)})=-\dfrac{4}{9}L^{-1}(\dfrac{1}{s+3})"




"+\dfrac{4}{9}L^{-1}(\dfrac{1}{s})-\dfrac{1}{3}L^{-1}(\dfrac{1}{s^2})"




"=-\dfrac{4}{9}e^{-3t}+\dfrac{4}{9}-\dfrac{1}{3}t"